# Scalable and robust regression models for continuous proportional data

### v1

@@ 5 Simulation studies
5.1 Robustness of cobin regression estimator
We conducted a simulation study to empirically validate Proposition 1, focusing on the consistency and robustness of the cobin and beta regression estimators. We considered four different scenarios for the response distribution: (A) beta, (B) cobin, (C) mixture of beta and uniform distributions, called the beta rectangular distribution (Bayes et al.,, 2012), and (D) mixture of three beta distributions. Specifically, the densities of (C) and (D) with mean $\mu$ correspond to
$p_{\mathrm{brec}}(y;\mu,\alpha,\phi)$ $=w(\mu,\alpha)p_{\mathrm{beta}}\left(y;\frac{\mu-0.5+0.5w(\mu,\alpha)}{w(\mu,\alpha)},\phi\right)+1-w(\mu,\alpha),$ (14)
$p_{\mathrm{bmix}}(y;\mu,\phi)$ $=0.25p_{\mathrm{beta}}(y;\mu-\epsilon(\mu),\phi)+0.5p_{\mathrm{beta}}(y;\mu,\phi)+0.25p_{\mathrm{beta}}(y;\mu+\epsilon(\mu),\phi),$ (15)
for $y\in(0,1)$, where $p_{\mathrm{beta}}(y;\mu,\phi)$ corresponds to the beta density parametrized by mean $\mu$ and precision $\phi$ (sum of two beta shape parameters), $w(\mu,\alpha)=1-\alpha(1-|2\mu-1|)$ for some $\alpha\in(0,1)$, and $\epsilon(\mu)=\min(\mu,1-\mu)/2$. We also consider two different link functions (cobit, logit) with sample sizes $n\in\{100,400,1600\}$, resulting in $4\times 2\times 3=24$ different data generating scenarios. We consider two covariates including the intercept and set the true coefficients as $\bm{\beta}=(\beta_{0},\beta_{1})=(0,1)$. The non-intercept covariate is generated as $x_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}N(0,\sigma_{x}^{2})$, where we set $\sigma_{x}^{2}=9$ for cobit and $\sigma_{x}^{2}=1$ for logit to account for the difference in scales between the link functions reflecting that $g^{\prime}_{\mathrm{cobit}}(0.5)$ and $g^{\prime}_{\mathrm{logit}}(0.5)$ differ by a factor of 3. For parameters not related to the mean, we set $\phi=8$ for beta, $\lambda=3$ for cobin, $(\alpha,\phi)=(0.2,10)$ for beta rectangular and $\phi=40$ for mixture of beta. For the beta rectangular, the choice of $\alpha=0.2$ ensures that the weight $w(\mu,\alpha)$ in (14) assigned to the beta distribution is greater than 0.8. We simulate 1000 replicated datasets for each data generation scenario. For each data set, we fit the cobin and beta regression models to find $\hat{\bm{\beta}}$ with unknown dispersion parameters, using the correct link function but not necessarily the true data generating distribution. We used iteratively reweighted least squares to find $\hat{\bm{\beta}}$ for cobin regression based on the stats package in R (R Core Team,, 2024), and we utilized the betareg R package (Cribari-Neto and Zeileis,, 2010) for the beta regression estimate.
Table 1: Analysis of beta and cobin regression estimates, better results are highlighted in bold. Results are based on 1000 replicates under the correct mean structure but possibly misspecified response distributions.
Beta Cobin Beta rectangular Mixture of beta
@@ 5.2 Robustness and scalability under spatial regression models
Next, we consider a spatial regression simulation study to analyze the robustness of cobin and micobin regression models against model misspecification, and highlight the computational benefits of the Kolmogorov-Gamma augmentation, particularly for latent Gaussian models. For data generation, we choose spatial locations uniformly at random from $[0,1]^{2}$ as training and test locations with sizes $(n_{\mathrm{train}},n_{\mathrm{test}})\in\{(200,50),(400,100)\}$. Then, spatial random effects are generated from a mean zero Gaussian process (GP) with exponential kernel $\operatorname*{cov}\{u(s),u(s^{\prime})\}=\sigma_{u}^{2}\exp(-\|s-s^{\prime}\|_{2}/\rho)$, where we set $\sigma_{u}^{2}=1$ and $\rho\in\{0.1,0.2\}$. For the fixed effect terms, we consider two covariates including the intercept, generate a non-intercept covariate as $x(s_{i})\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}N(0,3^{2})$, and set true coefficients as $\bm{\beta}=(\beta_{0},\beta_{1})=(0,1)$. The responses $y(s_{i})$ are generated from a beta rectangular distribution (14) with cobit link $\mu(s_{i})=g_{\mathrm{cobit}}^{-1}\{x(s_{i})\beta_{1}+u(s_{i})\}$, $\alpha=0.2$, and $\phi=10$. The beta rectangular response distribution leads to data that are occasionally sampled from the uniform distribution, reflecting potential outliers often encountered in realistic scenarios. This data generation process is repeated 100 times.
Based on the training set with size $n_{\mathrm{train}}$, we fit three different spatial regression models (beta, cobin, micobin) with correct mean structure (7) but none of the response distributions are correctly specified. We set a normal prior on the regression coefficients $\bm{\beta}\sim N_{2}(\bm{0}_{2},100^{2}\mathbf{I}_{2})$ and a half-Cauchy prior on the spatial random effects standard deviation $\sigma_{u}$. The spatial range parameter $\rho\in\{0.1,0.2\}$ is fixed at the true value. For spatial cobin and micobin regression models, we employ a blocked Gibbs sampler with Kolmogorov-Gamma augmentation. For the spatial beta regression model, we use Stan to carry out posterior inference using the No-U-Turn Sampler algorithm (Carpenter et al.,, 2017). We run a total of 6,000 MCMC iterations and record wall-clock running time, with the first 1,000 samples discarded as burn-in. For further details of the simulation settings, including priors for dispersion parameters, see Appendix S.5.
Table 2: Spatial regression simulation results under misspecified response distribution based on 100 replicates. Better results are highlighted in bold except for sampling performance.
Inference $(\hat{\beta}_{1})$ Prediction Sampling $(\bm{\beta})$
$\rho$ Method $(n_{\mathrm{train}},n_{\mathrm{test}})$ Bias RMSE negtestLL MSPE×$\times 10^{2}$ mESS time (min)
beta regression $(200,50)$0.048 0.118 -0.325 0.427 919.8 44.5
$(400,100)$0.052 0.089 -0.354 0.345 978.7 437.7
0.1 cobin regression $(200,50)$ 0.005 0.093 -0.340 0.388 2791.3 2.0
@@
beta regression $(200,50)$0.065 0.120 -0.320 0.329 1187.2 96.3
$(400,100)$0.052 0.095 -0.350 0.248 808.0 933.4
0.2 cobin regression $(200,50)$ 0.000 0.088 -0.346 0.306 3366.0 2.2
@@
Monte Carlo standard errors are all less than 0.015 for negtestLL, 0.013 for MSPE, 127.2 for mESS.
@@
Table 2 summarizes the results. First, with respect to the posterior mean estimate $\hat{\beta}_{1}$ of the fixed effect coefficient, cobin regression consistently produces the lowest bias and RMSE in all scenarios, in accordance with the previous simulation results. In contrast, beta and micobin regression estimates exhibit bias, with the magnitude of these biases increasing as spatial dependence becomes stronger. Second, to evaluate predictive performance under model misspecification, we compare the negative test log-likelihood conditional on random effects (negtestLL) and mean square prediction error (MSPE) based on test data. Micobin regression outperforms the others, while beta regression performs the worst on both metrics, highlighting the robustness of micobin regression in prediction. Finally, in terms of computational efficiency, cobin and micobin achieve a significantly higher multivariate effective sample size (mESS) (Vats et al.,, 2019) of $\bm{\beta}$ per unit time compared to beta regression. These findings demonstrate the scalability and robustness of cobin and micobin regression in estimation and prediction, respectively.
@@
As an illustrative application of cobin and micobin regression models with random effects, we analyze the benthic macroinvertebrate multimetric index (MMI) of US lakes and the association with lake watershed covariates. MMI, also known as an index of biotic integrity, is a standard quantitative measure for the bioassessment of macroinvertebrate assemblages (Karr,, 1991; Stoddard et al.,, 2008) that integrates various attributes of the assemblage (e.g. taxonomic composition and richness). Higher MMI values indicate a healthier and more diverse benthic macroinvertebrate community. We consider MMI data from the 2017 National Lake Assessment Survey (NLA), which covers about 1,000 lakes in the conterminous US; see the left panel of Figure 1. We refer to the 2017 NLA survey technical report for details on MMI (U.S. Environmental Protection Agency,, 2022).
@@
We are interested in understanding how the biotic integrity of lake ecosystems measured by MMI is associated with natural and human-related lake watershed characteristics, as well as in predicting the MMI of unsurveyed lakes. We consider LakeCat data (Hill et al.,, 2018) covering more than 380,000 US lakes, containing various natural and anthropogenic watershed covariates. For illustrative purposes, we select 7 watershed covariates that are highly important in the analysis of lake eutrophication (Hill et al.,, 2018, Fig. 7), as well as 2 additional covariates (manure application and urban land cover); see Table S.3 in Appendix S.6 for description. Since all covariates were heavily right-skewed and exhibited limited variation around the mean, we applied the transformation $x\mapsto\log_{2}(x+1)$ to reduce skewness and mitigate the influence of very large values.
Among the 2017 NLA survey data, a total of 950 lakes remained after removing those missing MMI and/or covariates. One lake had an MMI value of 0, and we removed it to allow a comparison with spatial beta and cobin regression models; the spatial micobin regression result with original data is available in Appendix S.6.

### v2

@@ 4 Simulation studies
4.1 Robustness of MLE to model misspecification
The first simulation study aims to empirically validate Proposition 1, focusing on the consistency and robustness of the maximum likelihood estimate of cobin regression models. We considered four different scenarios for the response distribution: (A) beta, (B) cobin, (C) mixture of beta and uniform distributions, known as the beta rectangular distribution (Hahn, , 2008), and (D) mixture of three beta distributions. Specifically, the densities of (C) and (D) with mean $\mu$ correspond to
$p_{\mathrm{brec}}(y;\mu,\alpha,\phi)$ $=w(\mu,\alpha)p_{\mathrm{beta}}\left(y;(\mu-0.5)/w(\mu,\alpha)+0.5,\phi\right)+1-w(\mu,\alpha),$ (12)
$p_{\mathrm{bmix}}(y;\mu,\phi)$ $=0.25p_{\mathrm{beta}}(y;\mu-\epsilon(\mu),\phi)+0.5p_{\mathrm{beta}}(y;\mu,\phi)+0.25p_{\mathrm{beta}}(y;\mu+\epsilon(\mu),\phi),$
for $y\in(0,1)$, $w(\mu,\alpha)=1-\alpha(1-|2\mu-1|)$ for some $\alpha\in(0,1)$, and $\epsilon(\mu)=\min(\mu,1-\mu)/2$. We also consider two different link functions (cobit, logit) with sample sizes $n\in\{100,400,1600\}$, resulting in $4\times 2\times 3=24$ different data generating scenarios. We consider two covariates including the intercept and set the true coefficients as $\bm{\beta}=(\beta_{0},\beta_{1})=(0,1)$. The non-intercept covariate is generated as $x_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}N(0,\sigma_{x}^{2})$, where we set $\sigma_{x}^{2}=9$ for cobit and $\sigma_{x}^{2}=1$ for logit to account for the difference in scales between the link functions reflecting that $g^{\prime}_{\mathrm{cobit}}(0.5)$ and $g^{\prime}_{\mathrm{logit}}(0.5)$ differ by a factor of 3. For parameters not related to the mean, we set $\phi=8$ for beta, $\lambda=3$ for cobin, $(\alpha,\phi)=(0.2,10)$ for beta rectangular and $\phi=40$ for mixture of beta. For the beta rectangular, the choice of $\alpha=0.2$ ensures that the weight $w(\mu,\alpha)$ in (12) assigned to the beta distribution is greater than 0.8. We simulate 1000 replicated datasets for each data generation scenario.
For each data set, we fit the cobin and beta regression models to find MLE $\hat{\bm{\beta}}$ with unknown dispersion parameters, using the correct link function but not necessarily the true data generating distribution. We used iteratively reweighted least squares to find $\hat{\bm{\beta}}$ for cobin regression based on the stats package in R (R Core Team, , 2024), and we utilized the betareg R package (Cribari-Neto and Zeileis, , 2010) for the beta regression estimate.
Table 1: Analysis of beta and cobin regression MLE under four different data generating distributions: Beta, Cobin, beta rectangular (Betarec), and mixture of beta (Beta3mix). Results are based on 1000 replicates under the correct mean structure but possibly misspecified distributions.
Beta data Cobin data Betarec data Beta3mix data
@@ 4.1 Robustness of MLE to model misspecification
Supplementary Section S.5.1 presents additional results, including a visualization of the data-generating distributions, the length of 95% confidence intervals and empirical coverage probabilities, and predictive performance under even misspecified mean structures. These results further demonstrate that cobin regression is more robust than beta regression in terms of both inference and prediction across a range of misspecified settings.
@@ 4.2 Robustness and scalability under spatial regression models
We consider a spatial regression simulation study to analyze the robustness of cobin and micobin regression models to distributional misspecification. This simulation also aims to assess the scalability achieved through the Kolmogorov-Gamma augmentation, particularly for latent Gaussian hierarchical models. For data generation, we choose spatial locations uniformly at random from $0,1]^{2}$ as training and test locations with sizes $(n_{\mathrm{train}},n_{\mathrm{test}})\in\{(200,50),(400,100)\}$. Then, spatial random effects are generated from a mean zero Gaussian process (GP) with exponential kernel $\operatorname*{cov}\{u(s),u(s^{\prime})\}=\sigma_{u}^{2}\exp(-\|s-s^{\prime}\|_{2}/\rho)$, where we set $\sigma_{u}^{2}=1$ and $\rho\in\{0.1,0.2\}$. For the fixed effect terms, we consider two covariates including the intercept, generate a non-intercept covariate as $x(s_{i})\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}N(0,3^{2})$, and set true coefficients as $(\beta_{0}^{\mathrm{true}},\beta_{1}^{\mathrm{true}})=(0,1)^{\mathrm{\scriptscriptstyle T}}$. The responses $y(s_{i})$ are generated from a beta rectangular distribution (12) with cobit link $\mu(s_{i})=g_{\mathrm{cobit}}^{-1}\{\beta_{0}^{\mathrm{true}}+x(s_{i})\beta_{1}^{\mathrm{true}}+u(s_{i})\}$, $\alpha=0.2$, and $\phi=10$. This data generation process is repeated 100 times.
Based on the training set with size $n_{\mathrm{train}}$, we fit three different spatial regression models (beta, cobin, micobin) with correct mean structure (8) but none of the response distributions are correctly specified. We set a $N_{2}(\bm{0}_{2},100^{2}\mathbf{I}_{2})$ prior on the regression coefficients $\bm{\beta}$ and a half-Cauchy prior on the spatial random effects standard deviation $p(\sigma_{u})\propto(1+\sigma_{u}^{2})^{-1}$, $\sigma_{u}>0$. The spatial range parameter $\rho\in\{0.1,0.2\}$ is fixed at the true value. For spatial cobin and micobin regression models, we employ a blocked Gibbs sampler with Kolmogorov-Gamma augmentation. For the spatial beta regression model, we use Stan to carry out posterior inference using the No-U-Turn Sampler algorithm (Carpenter et al., , 2017). We run a total of 6,000 MCMC iterations and record wall-clock running time, with the first 1,000 samples discarded as burn-in.
Table 2: Bayesian spatial regression simulation results based on 100 replicates, under the correct mean structure but misspecified distributions.
Spatial dependence Inference $(\hat{\beta}_{1})$ Prediction Sampling $(\bm{\beta})$
[…]
beta regression $(200,50)$ -0.048 0.118 -0.325 0.427 919.8 44.5
$(400,100)$ -0.052 0.089 -0.354 0.345 978.7 437.7
$\rho=0.1$ (moderate) cobin regression $(200,50)$ 0.005 0.093 -0.340 0.388 2791.3 2.0
@@
beta regression $(200,50)$ -0.065 0.120 -0.320 0.329 1187.2 96.3
$(400,100)$ -0.052 0.095 -0.350 0.248 808.0 933.4
$\rho=0.2$ (strong) cobin regression $(200,50)$ 0.000 0.088 -0.346 0.306 3366.0 2.2
@@
NTLL, negative test log-likelihood; MSPE, mean square prediction error; mESS, multivariate effective sample size. Monte Carlo standard errors are all less than 0.015 for NTLL, 0.013 for MSPE, 127.2 for mESS.
@@
Table 2 summarizes the results. First, with respect to the posterior mean estimate $\hat{\beta}_{1}$ of the fixed effect coefficient, cobin regression consistently produces the lowest bias and RMSE in all scenarios, in accordance with the previous simulation results. In contrast, beta and micobin regression estimates exhibit bias, with the magnitude of these biases increasing as spatial dependence becomes stronger, which is partly attributable to beta and micobin not belonging to the exponential dispersion model. Second, to evaluate predictive performance under model misspecification, we compare the negative test log-likelihood conditional on random effects (negtestLL) and mean square prediction error (MSPE) based on test data. Micobin regression outperforms the others, while beta regression performs the worst on both metrics, highlighting the robustness of micobin regression in prediction. Finally, in terms of computational efficiency, cobin and micobin achieve a significantly higher multivariate effective sample size (mESS) (Vats et al., , 2019) of $\bm{\beta}$ per unit time compared to beta regression. These results highlight the scalability of both cobin and micobin regressions in hierarchical settings, with cobin showing more reliable parameter estimation, consistent with the behavior suggested by Proposition 1 in a simpler setting, and micobin offering greater robustness in predictive performance under potential distributional misspecification. Supplementary Section S.5.2 contains further details of the spatial regression simulation settings, including priors for dispersion parameters, as well as additional summaries including 95% credible interval length and empirical coverage probabilities of $\beta_{1}$.
4.3 Robustness to the presence of outliers
We conduct a third simulation study to systematically assess the robustness of the cobin and micobin models to outliers, focusing on the stability of both parameter estimates and predictions. We considered two types of datasets: one without outliers, denoted as $\mathcal{D}=\{(y_{i},\bm{x}_{i})\}_{i=1}^{n}$ with size $n=500$, and one with a single outlier, denoted as $\mathcal{D}^{\circ}=\mathcal{D}\cup\{(y^{\circ},\bm{x}^{\circ})\}$ with size $501$. We set $p=3$ predictors including an intercept, where covariates in $\mathcal{D}$ are generated as $(x_{i1},x_{i2})\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}N_{2}(\bm{0},3^{2}\mathbf{I}_{2})$, $i=1,\dots,n$. We considered two different response distributions for $\mathcal{D}$, (A) beta and (B) cobin, under the same mean structure $g_{\mathrm{cobit}}\{E(y_{i}\mid\bm{x}_{i})\}=\beta_{0}+\beta_{1}x_{i1}+\beta_{2}x_{i2}$. The true regression coefficients were set as $(\beta_{0}^{\mathrm{true}},\beta_{1}^{\mathrm{true}},\beta_{2}^{\mathrm{true}})^{\mathrm{\scriptscriptstyle T}}=(-6,1,0)^{\mathrm{\scriptscriptstyle T}}$, and for dispersion parameters we set $\phi=17$ for beta and $\lambda=6$ for cobin. We then introduced an outlier with covariate $\bm{x}^{\circ}=(6,6)^{\mathrm{\scriptscriptstyle T}}$ and a response value of either $y^{\circ}=10^{-2}$ (setting 1) or $y^{\circ}=10^{-3}$ (setting 2), resulting in four different data settings (A1, A2, B1, B2). Since $E(y\mid\bm{x}=\bm{x}^{\circ})=0.5$ under the true model without outlier, the response $y^{\circ}=10^{-2}$ or $y^{\circ}=10^{-3}$ are highly unlikely, with $y^{\circ}=10^{-3}$ being more extreme. This data generation process was repeated 100 times.
For each setting with and without outliers, we fit four different fixed-effects models (beta, cobin, beta rectangular, micobin) with the mean structure $g_{\mathrm{cobit}}\{E(y_{i}\mid\bm{x}_{i})\}=\beta_{0}+\beta_{1}x_{i1}+\beta_{2}x_{i2}$ for $i=1,\dots,n$. We denote $\hat{\beta}_{j}^{\circ}$ and $\hat{\beta}_{j}$ $(j=0,1,2)$ as posterior mean estimates of $\beta_{j}$ based on the dataset with outlier ($\mathcal{D}^{\circ}$) and without outlier $(\mathcal{D})$. For all four models, we set a $N_{3}(\bm{0}_{3},100^{2}\mathbf{I}_{3})$ prior on the regression coefficients $\bm{\beta}$. For the cobin and micobin regression models, posterior computation is performed with Kolmogorov-Gamma augmentation based on Algorithm 1 and 2. For beta regression and beta rectangular regression, we use Stan with the No-U-Turn Sampler. We ran a total of 6,000 MCMC iterations and discarded the first 1,000 samples as a burn-in.
Table 3: Outlier robustness simulation results based on 100 replicates in terms of stability of parameter estimates $|\Delta\hat{\beta}_{j}|$, empirical coverage, and stability of linear predictor $\|\Delta\hat{\eta}_{1:n}\|_{2}$.
Method Setting $|\Delta\hat{\beta}_{0}|$ $|\Delta\hat{\beta}_{1}|$ Coverage of $\mathrm{CI}_{.95}(\beta_{1})$ $|\Delta\hat{\beta}_{2}|$ Coverage of $\mathrm{CI}_{.95}(\beta_{2})$ $\|\Delta\hat{\eta}_{1:n}\|_{2}$
with $y^{\circ}$ without $y^{\circ}$ with $y^{\circ}$ without $y^{\circ}$
beta regression A1 0.095 0.061 74.0% 94.0% 0.056 69.0% 96.0% 5.948
A2 0.145 0.091 51.0% 0.083 39.0% 8.852
cobin regression A1 0.007 0.017 95.0% 97.0% 0.020 95.0% 95.0% 1.768
A2 0.008 0.016 95.0% 0.020 96.0% 1.762
betarec regression A1 0.039 0.007 93.0% 92.0% 0.001 95.0% 96.0% 0.987
A2 0.038 0.006 93.0% 0.001 96.0% 0.964
micobin regression A1 0.007 0.005 91.0% 90.0% 0.006 93.0% 95.0% 0.601
A2 0.006 0.005 92.0% 0.006 93.0% 0.562
beta regression B1 0.099 0.062 41.0% 88.0% 0.058 54.0% 92.0% 6.136
B2 0.150 0.093 17.0% 0.085 28.0% 9.111
cobin regression B1 0.004 0.017 96.0% 96.0% 0.020 91.0% 96.0% 1.770
B2 0.005 0.017 97.0% 0.021 93.0% 1.807
betarec regression B1 0.044 0.007 77.0% 79.0% 0.001 91.0% 91.0% 1.114
B2 0.045 0.007 75.0% 0.001 91.0% 1.124
micobin regression B1 0.005 0.003 93.0% 91.0% 0.004 94.0% 96.0% 0.357
B2 0.005 0.003 93.0% 0.004 95.0% 0.368
Monte Carlo standard errors are all less than 0.003 for $|\Delta\hat{\beta}_{j}|$ ($j=0,1,2$), 5.0% for coverage probabilities, and 0.098 for $\|\Delta\hat{\eta}_{1:n}\|_{2}$.
We summarize simulation results in three different ways. First, we compute $|\Delta\hat{\beta}_{j}|=|\hat{\beta}_{j}^{\circ}-\hat{\beta}_{j}|$ to quantify the stability of the parameter estimates in the presence of an outlier. Also, we compare the empirical coverage of the 95% credible intervals for $\beta_{1}^{\circ},\beta_{1}$ (with $\beta_{1}^{\mathrm{true}}=1$) and $\beta_{2}^{\circ},\beta_{2}$ (with $\beta_{2}^{\mathrm{true}}=0$) to assess how outliers affect both point estimates and uncertainty quantification. Finally, we assess the stability of the linear predictor using $\|\Delta\hat{\eta}_{1:n}\|_{2}=\left[\sum_{i=1}^{n}(\hat{\eta}_{i}^{\circ}-\hat{\eta}_{i})^{2}\right]^{1/2}$, where $\hat{\eta}_{i}=\hat{\beta}_{0}+\hat{\beta}_{1}x_{i1}+\hat{\beta}_{2}x_{i2}$ and $\hat{\eta}_{i}^{\circ}=\hat{\beta}_{0}^{\circ}+\hat{\beta}_{1}^{\circ}x_{i1}+\hat{\beta}_{2}^{\circ}x_{i2}$.
Table 3 demonstrates that beta regression is the most sensitive to outliers in all aspects. In terms of stability of regression coefficients and linear predictors, beta regression exhibits poor performance, especially showing a stark difference between settings where $y^{\circ}=10^{-2}$ and $y^{\circ}=10^{-3}$. This sharp contrast is also reflected in the empirical coverage; all coverage values fall below 74% in Setting 1 ($y^{\circ}=10^{-2}$) and below 51% in Setting 2 ($y^{\circ}=10^{-3}$). In contrast, the cobin regression result is virtually unaffected by outliers. Both beta rectangular and micobin regressions, which are based on mixture formulations, offer generally better stability than cobin regression. Between the two, micobin regression shows greater robustness in terms of the stability of the linear predictor. This is partly because the intercept estimate in beta rectangular regression is relatively non-robust (see $|\Delta\hat{\beta}_{0}|$ column), since the beta rectangular model assumes outliers arise from a uniform distribution with fixed mean 1/2, which tends to highly influence the intercept estimate.
Overall, cobin regression provides clear advantages over beta regression in terms of robustness to outliers, even without relying on mixture formulations. Micobin regression achieves even greater robustness, with similar improvements. However, the beta rectangular model suffers from instability in the intercept estimate, which leads to degraded predictive performance compared to the micobin regression model. Supplementary Section S.5.3 presents additional results, including visualizations of the data-generating distributions, prior distributions for parameters not related to the mean, and summaries of posterior mean estimates $\hat{\beta}_{j}^{\circ}$, $\hat{\beta}_{j}$ and sampling efficiency. These additional results further support the robustness of cobin and micobin regression methods in the presence of outliers.
@@ 4.3 Robustness to the presence of outliers
As an illustrative application of cobin and micobin regression models, we analyze the benthic macroinvertebrate multimetric index (MMI) of US lakes and the association with lake watershed covariates. MMI, also known as an index of biotic integrity, is a standard quantitative measure for the bioassessment of macroinvertebrate assemblages (Karr, , 1991; Stoddard et al., , 2008) that integrates various attributes of the assemblage (e.g. taxonomic composition and richness). Higher MMI values indicate a healthier and more diverse benthic macroinvertebrate community. We consider MMI data from the 2017 National Lake Assessment Survey (NLA) (U.S. Environmental Protection Agency, , 2022), which covers about 1,000 lakes in the conterminous US; see the left panel of Figure 1.
@@
We are interested in understanding how the biotic integrity of lake ecosystems measured by MMI is associated with natural and human-related lake watershed characteristics, as well as in predicting the MMI of unsurveyed lakes. We consider LakeCat data (Hill et al., , 2018) covering more than 380,000 US lakes, containing various natural and anthropogenic watershed covariates. For illustrative purposes, we select 7 watershed covariates that are highly important in the analysis of lake eutrophication (Hill et al., , 2018, Fig. 7), as well as 2 additional covariates (manure application and urban land cover); see Table S.6 in Supplementary Section S.6.1 for description. Since all covariates were heavily right-skewed and exhibited limited variation around the mean, we applied the transformation $x\mapsto\log_{2}(x+1)$ to reduce skewness and mitigate the influence of very large values.
Table 4: Lakes with the 3 lowest MMI values (multiplied by 0.01 from the original 0–100 scale)
COMID Longitude Latitude Lake name, State MMI Mean MMI (10 nearest) $\mathcal{D}^{\star}$ $\mathcal{D}^{\circ}$ $\mathcal{D}$
22721231 -81.893 33.300 Brierpatch Lake, GA 0 0.534
9201925 -80.263 35.116 Jones Pond, NC 0.020 0.460
22845861 -92.273 34.538 Ferguson Lake, AR 0.021 0.414
COMID: unique lake identifier; checkmark indicates inclusion in the respective dataset ($\mathcal{D}^{\star}$, $\mathcal{D}^{\circ}$, or $\mathcal{D}$).
After excluding lakes with missing MMI and/or covariates, a total of 950 lakes remained. Among 950 lakes, MMI values range from 0 to 0.904 (mean 0.452; median 0.450, Q1 0.340, Q3 0.564), with one lake at 0 and two near the lower boundary (0.020, 0.021). To assess robustness, we analyze three datasets: (1) the full dataset $(\mathcal{D}^{\star},n=950)$; (2) one lake with zero MMI removed $(\mathcal{D}^{\circ},n=949)$; and (3) three lakes removed $(\mathcal{D},n=947)$, additionally excluding the two low MMI lakes. As shown in Table 4, these lakes exhibit exceptionally low MMI compared to their surroundings, potentially due to localized environmental degradation such as waste discharge or nearby construction activity. In fact, beta and cobin regression fit results on $\mathcal{D}^{\circ}$ reveal that the two low MMI lakes (0.020 and 0.021) exhibit poor goodness-of-fit based on residual diagnostics, suggesting they are acting as potential outliers under these models (see Supplementary Section S.6.2 for details). By comparing the analysis results with and without these potential outliers, we aim to evaluate the robustness of the proposed cobin and micobin regression models and compare them with beta regression.